OpenStudy (anonymous):

Given f(x)= {5x-6, x<2 {x^2, 24 find f'(x) and give its domain My question is how does taking the derivative affect the restrictions?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@zepdrix @jagr2713

OpenStudy (anonymous):

@freckles @dan815

OpenStudy (freckles):

the endpoints need to be check to see if left derivative=right derivative

OpenStudy (freckles):

for example for f' to exist at x=2 you must have (5x-6)'=(x^2)' as x goes to 2

OpenStudy (freckles):

you will need to do something similar for x=4

OpenStudy (freckles):

oh you closed this does that mean you figured it out?

OpenStudy (anonymous):

no i didn't figure it out

OpenStudy (anonymous):

ill try to take a screenshot of the problem

OpenStudy (freckles):

are you understanding what I'm saying though?

OpenStudy (freckles):

f' is definitely going to exist for all x except for x=2 or x=4 because all the functions are continuous for all real numbers for x=2 or x=4 we are not sure in this case we have to check to see if the left derivative=right derivative for both x=2 and x=4

OpenStudy (anonymous):

so its a limit problem? i need to look at whether theres a limit at 2 and 4?

OpenStudy (freckles):

\[\lim_{x \rightarrow 2}(5x-6)'=\lim_{x \rightarrow 2}(x^2)' \text{ if this holds then yes } \\ f' \text{ exists at } x=2 \] \[\lim_{x \rightarrow 4}(x^2)'=\lim_{x \rightarrow 4}(2x+8)' \text{ if this holds then } f' \text{ exists at } x=4\] you are checking to see if the derivative is continuous at these numbers we are doing exactly what would do to see if the original function was continuous at x=2 or x=4 which it is since 5*2-6=2^2 and 4^2=2*4+8 which should have been the first thing we looked at but anyways back to the above limits

OpenStudy (anonymous):

so the first one does exist because its 4=4 and the second one exists too because its 16=16?

OpenStudy (anonymous):

so the answer is B, domain= all real numbers?

OpenStudy (freckles):

how did you get that?

OpenStudy (anonymous):

i plugged in for x for each equation

OpenStudy (freckles):

after finding the derivative right?

OpenStudy (freckles):

I already checked above to see if the original function was continuous at x=2 or x=4 again we need to check to see if the derivative function is continuous at x=2 or x=4

OpenStudy (anonymous):

the derivative of the first would equal 5 and of the second 2x?

OpenStudy (freckles):

\[\lim_{x \rightarrow 2}(5x-6)'=\lim_{x \rightarrow 2}(x^2)' \\ \lim_{x \rightarrow 2}(5)=\lim_{x \rightarrow 2}(2x)\] ok now see if this equality holds

OpenStudy (anonymous):

no because it would be 5=4

OpenStudy (freckles):

right so f' will definitely not exist at x=2 since the slopes do not match up at x=2

OpenStudy (anonymous):

so it would exist at 4 either because the derivative of the first is 2x and 2 for the second. so it would be 8=2

OpenStudy (anonymous):

so the answer is A

OpenStudy (freckles):

right To find if f'(a) exists: see if f is continuous at a and if yes continue to see if f' is continuous at a

OpenStudy (anonymous):

ok can you help with one more please?

OpenStudy (freckles):

might be able to

OpenStudy (freckles):

and you graphed it?

OpenStudy (anonymous):

yes, im not sure what it means increases without bound

OpenStudy (anonymous):

 it looks like this

OpenStudy (freckles):

it means it function increases and never stops

OpenStudy (anonymous):

but wouldnt there also be a tangent line of 0 at the maximum?

OpenStudy (freckles):

what kind of tangent line ? :)

OpenStudy (anonymous):

 wouldn't that be a tangent line? the tangent line is the slope of a derivative?